# Cracking Vector Search Indexes

### v1

@@ 3.3. Outline of the solution
Based on these insights, we propose a partition-based crackable IVF index for ANN search. It is designed to remain near the optimal minimum cumulative time across the number of queries submitted (Figure 1), improving as the workload increases, and addressing the core issue of indexing parts of the search space that are not used.
CrackIVF employs two localized and independent build operations: CRACK and REFINE. The former introduces new partitions, while the latter applies a localized K-means to improve them. These operations run independently across both time and vector space.
We directly follow the query distribution, where queries themselves act as candidates for CRACK and REFINE, determining where new partitions should be introduced, so they are not randomly sampled, and which local regions need refinement. Intermediate results from search operations, such as the distances computed to nearby points, can be re-used to amortize the cost of assignments. Using the queries to create partitions, rather than selecting random ones, allows distances computed during search to reused, steal points from the local region, that are closer to the query than their current assignment. Additionally, using the queries as candidates for CRACK and REFINE enables CrackIVF to exactly follow the query distribution in terms of when and where to execute build operations, as partitions that receive a larger number of queries will proportionally be candidates for cracking and refinement by the same amount.
Finally, to mitigate the high overhead of storage reorganization in high-dimensional IVF indexes, we do not perform build operations after every query. Instead, CRACK operations are buffered and committed when beneficial, while REFINE is executed eagerly but infrequently to improve local regions that queries are visiting when necessary. To automate these decisions, we introduce two control mechanisms. The first is a set of heuristic rules, that determine where in the vector space cracks should be buffered or refinements executed. The second is a cost-based approach that estimates the latency overhead of a build operation and constrains it relative to the cumulative runtime of the index since initialization. As more queries arrive and the index is used more frequently, the budget for build operations increases proportionally, allowing the index to grow and refine in sync with the query workload.
@@
[…]
@@
$C_{\text{local}}\in\mathbb{Z}^{nprobe}$ Local region centroid IDs
@@
$\max_{\text{points}}$ K-means max training sample per centroid
$nlist$ Total inverted lists storing $P$, where $nlist=|C|$
@@
$bs$ Total queries batched in single SEARCH
$nprobe$ Number of partitions scanned per SEARCH
@@
$A_{P}^{\ast}\in\mathbb{Z}^{|P|}$ Point assignments to centroid IDs
$D_{P}^{\ast}\in\mathbb{R}^{|P|}$ Point distances to centroids
$H_{C}^{\ast}\in\mathbb{N}^{|C|}$ Histogram of points per centroid
[…]
Table 1. Notation used in algorithms and cost model.
@@ 4. Crack-IVF
CrackIVF, is a partition-based crackable IVF index implemented in FAISS (Douze et al., 2024). The core logic of the algorithm and operations are covered in Section 4.1 and Algorithm 1. Control mechanisms, including heuristic rules and cost models, are described in Section 4.2 and Section 4.3. They aid in making accurate decisions about where and when to perform the build operations, CRACK and REFINE, which are shown in Algorithm 2 and Algorithm 3. The notation used in this section is summarized in Table 1.
@@ 4.1. Algorithm and Operations
CrackIVF performs index construction and storage reorganization operations as a side-effect of query execution to incrementally improve index performance. It has three core operations, $SEARCH$ finds the k-NN results to return to the user, and $CRACK$, $REFINE$ are the build operations that perform physical reorganization of the IVF structure as well as change the index state.
SEARCH: The search procedure is common among partition based IVF indexes, where the local region to the query is identified by the nearest partitions, whose point assignments are subsequently scanned to find the final $k$ nearest neighbors of the query. In our case, the result are point ids, $I$, and distances $D$ to the query, of the visited points in the local region sorted by similarity, together with the ids of the cracks that the query visited, $C_{visited}$, as seen in Line 7 of Algorithm 1.
@@
8:     $I_{\text{steal}}\leftarrow\{p\in I_{\text{local}}\mid D_{\text{local}}(p)\mathrel{\text{op}}D_{P}^{\text{dyn}}(p)\}$op = $<$ for L2, $>$ for IP}
9:     $good\_crack\_candidate$ = CrackHeuristic($I_{steal},\text{State}_{\text{dyn}}$)
10:     if $good\_crack\_candidate$ then
@@
20:  if $can\_afford\_crack$ and $enough\_buffered$ then
@@
26:     CRACK$(C_{\text{buffered}},I_{\text{buffered}})$Algorithm 2}
@@
34:        if $can\_afford\_refine$ then
@@
36:           if $good\_refine\_candidate$ then
@@
CRACK: Physical reorganization of points between inverted lists, and the mutation of the index does not happen after each query. Each incoming query represents a crack candidate, which is evaluated by heuristic rules to classify it as good or bad (Section 4.2) (Line 9 Algorithm 1). Only good cracks are further buffered and potentially committed at a later time. Each crack is essentially a new partition, where the points assigned to the partition have been “stolen” based on their proximity to the query. We reuse the results of the SEARCH operations to compare if the distance of a point to the query is closer than the distance of the point to its current assignment, which could be another buffered crack. If it is, then it’s beneficial to reassign the point to the new crack, and this assignment is also buffered, to be committed when $CRACK$ is executed, as explained in Section 4.3. $CRACK$ is the procedure performing the physical reorganization, syncing the buffered state of the index with the currently working state. The core individual kernels of the procedure are sequentially executed (Algorithm 2). It adds a new inverted list for each buffered crack, moves stolen points to their new assignments, and updates the partition representatives to be the centroid of the points based on their new assignments. This updates the location of the crack representatives, whose point assignments were stolen, in order for their new location to match the underlying assignment distribution. This procedure is only executed when enough budget for build operations has been accumulated through the usage of the index (Line 18 Algorithm 1). This control mechanism ensures that CrackIVF does not spend a disproportionate amount of time on CRACK operations. The parameter $\alpha$, the ratio of total build operations time to total search time, is what controls the available budget. Thus, since $CRACK$ is lazy, the heuristics governing where a crack is added and the cost-based control mechanism for when it is added are independent.
@@
REFINE: Refinement is the second operation to improve the index. It is performed eagerly, focusing on the local region visited by the query. The goal is to immediately correct suboptimal regions in the current index. Thus, the buffered states and buffered cracks within the local region are not considered. The decision to refine a region is based on heuristic rules that evaluate the imbalance of point assignments in local cracks. Additionally, executing a refinement operation depends on an estimate of the runtime cost, determined using a cost model specific to $REFINE$ (Line 33 Algorithm 1).
Unlike cracking, the opportunity to refine a local region is transient. The heuristic rules must decide that the currently visited local region is a good candidate for refinement, and at the same time, there must be sufficient computational budget to execute the operation. If where and when to execute a refine do not simultaneously agree, the next opportunity for the same region will arise only when another query accesses it.
$REFINE$, as described in Algorithm 3, executes a sequence of kernels implementing a local K-means variation and performs physical reorganization. This process improves the placement of crack representatives within the local region and reassigns points to their nearest representative. Unlike $CRACK$, $REFINE$ incurs additional computational costs as it does not rely on precomputed buffered assignments. Instead, assignments are determined dynamically upon completion of the local K-means process.
@@
2:  Local K-Means: Refine region of $C_{\text{local}}$, producing $C^{\prime}_{\text{local}}$ centroids.
@@
Index State: Since CRACK is a lazy operation, the index can exist in two concurrent states, the true state and a dynamic state. The true state reflects the current structure of the index, while the dynamic one tracks its expected future structure for when $CRACK$ is executed, which synchronizes the two states. Maintaining both is essential. $REFINE$ operations directly alter the current index state, while cracking decisions are made based on the dynamic one, leading to scenarios that require us to undo buffered crack decisions. Initially, the heuristic rules considered some crack candidates as beneficial, and CrackIVF buffered them, but they have since been altered. Rolling back the dynamic state avoids bad cracks from being committed and having a situation where we commit decisions to the index that we know to be suboptimal (Line 21 Algorithm 1).
To track the dynamic, uncommitted, state of the index, we define:
(1) $\text{State}_{\text{dyn}}=(A_{P}^{\text{dyn}},D_{P}^{\text{dyn}},H_{C}^{\text{dyn}})$
which consists of dynamic assignments $A_{P}^{\text{dyn}}$, dynamic point distances $D_{P}^{\text{dyn}}$, and a histogram $H_{C}^{\text{dyn}}$ of true and buffered crack sizes.
Similarly, the true state is defined as:
(2) $\text{State}_{\text{true}}=(A_{P}^{\text{true}},D_{P}^{\text{true}},H_{C}^{\text{true}})$
representing the true, committed state of assignments, distances, and crack sizes in the index.
@@
The initialization of the above states happens during the index setup time. More specifically, in our FAISS-based implementation, during an “.add()“ operation. This is when the data vectors are assigned to their nearest partitions by calculating the distances to each partition representative. We extend this operation to return to the assignments and distances computed, which were an intermediate result. The state held in $A_{P}^{\text{dyn}}$,$D_{P}^{\text{dyn}}$ and $H_{C}^{\text{dyn}}$, is only consistent with the actual state of the index after $CRACK$ operations and before buffering any cracks. If no cracks are buffered and only $REFINE$ operations are executed, the dynamic and true state are always consistent. All of the above state metadata is used after the SEARCH operation, along with the search results $I,D,C_{local}$ to classify each query as a good or bad crack candidate (Section 4.2).
@@
Rollback of buffered state: Consider the following scenario. A $CRACK$ operation was just executed so at $t_{0}$ $State_{true}=State_{dyn}$. After a few queries, at $t_{1}$, assume that 2 good crack candidates have been buffered, both in the $C_{local}$ region of the vector space and thus now $State_{true}\neq State_{dyn}$. At a later point, $t_{2}$, a $REFINE$ is executed, also in $C_{local}$ refining it to $C^{\prime}_{local}$ and changing $S_{true}$. Since $REFINE$ operates to improve the current state eagerly, the decision that the buffered cracks are good was taken prior, at $t_{1}$ based on $C_{local}$, and thus they may now have had their points stolen back, and assigned to the refined cracks in $C^{\prime}_{local}$. Thus even though still $|C_{buffered}|=2$, it may now contain bad cracks. So when at $t_{3}$, it is a good time to execute a $CRACK$ operation again, all buffered cracks are re-evaluated, and only the good cracks that remain are committed, ensuring that in the end, we are back to $State_{true}=State_{dyn}$. This logic can be seen in lines 21-25 of Algorithm 1.
Memory Overhead and Optimizations: The additional memory cost of maintaining $State_{\text{dyn}}$ and $State_{\text{true}}$ is $O(4|P|+2|C|)$, which is negligible since $d\gg 4$, so in our in-memory implementation that data alone are $O(|P|*d)$. Additionally, these array structures are freed once the index converges or incoming queries stop. In our case, it enables an efficient compute-storage trade-off, deferring cracking costs instead of paying them per query.
@@
There are a number of optimizations that a fully fledged implementation can make use of. For embedding lakes, where vectors are stored on disk, state tracking arrays can be stored alongside data points instead of in the index (irrelevant for our in-memory FAISS implementation). Since queries for an individual index are typically infrequent, $State_{\text{dyn}}$ can be committed or flushed between queries, while $State_{\text{true}}$ is recomputed as needed. Full memory overhead is only necessary during bursty query loads, where cracking must be buffered to reduce cumulative overhead and prevent long tail response times. Finally, the cracking overhead can be hidden from query response times since results are available before any cracking operations begin. A separate thread pool can handle cracking operations while queries use the current index state, provided $CRACK$ and $REFINE$ remain atomic to prevent inconsistencies. Prior work on the related problem of index maintenance has demonstrated similar approaches to minimizing the overhead of mutating the ANN index state (Xu et al., 2023) with concurrent operations, we further cover these approaches in Section 6.
Our work assumes a continuous query load at the system’s capacity, leveraging inter-query parallelism. As a result, the memory overhead for index state tracking is unavoidable in our implementation if we aim to maintain high performance.
@@ 4.2. Where to apply build operations?
There is a constant stream of crack and refinement candidates, and as observed in Section 3, not all regions of the vector space are equally important, and not all operations applied lead to improvement of index performance. Thus, a set of rules is needed to distinguish good candidates from bad ones. The heuristic rules provided in this section aim to distill the learning from our initial observations to capture independent cases where we want to add a new crack or refine a region. At this stage, they are not an exhaustive list, as we plan to extend them in future work.
Where to CRACK: The CrackHeuristic determines whether an incoming query is a good or bad candidate for cracking, as shown in Algorithm 1. It evaluates the stolen point IDs $I_{\text{steal}}$ and the dynamic state $State_{\text{dyn}}$ to return a True/False decision. If classified as good, the query is buffered for a future $CRACK$.
By default, we aim to expand the index size, as our insights suggest that optimal performance is typically achieved at some maximum number of partitions. Since $REFINE$ does not add partitions, we assume every incoming query is a good candidate to have a constant stream of candidates that can expand the size of a CrackIVF index and only reject them if they pass any of our heuristic rules.
CrackHeuristic: A query is considered a bad candidate (returning False) if any of the following conditions hold:
Don’t Crack: Rule 1 rejects queries with too few stolen points:
@@
Don’t Crack: Rule 2 prevents excessive partitioning, constraining the number of partitions within a region relative to the number of points:
[…]
Where to REFINE: The RefineHeuristic determines whether a local region in the index is a good candidate for refinement, returning a $True/False$ value for if a refinement should take place. Unlike $CRACK$, where the latency overhead is amortized by buffering, $REFINE$ is applied eagerly and is also an expensive operation, running K-Means locally, calculating new point assignments, and reorganizing the physical storage. Additionally, if the index continues to grow in the future, any previously refined region may need further refinement, making premature refinements inefficient. It is much better to relax rules for when to refine when the index has reached its maximum size and has also accumulated a larger cost budget.
Therefore, we do not want to be calling $REFINE$ repeatedly, unless necessary, and our default behavior is to assume that every incoming query local region is NOT a good candidate for refinement unless it satisfies specific heuristic rules that indicate a significant local partition imbalance. Imbalance affects tail response time negatively and increases the variance since even though queries may be accessing the same number of partitions, they may be scanning a vastly different number of points. Additionally, running K-means improves the centroid location of the crack representatives, making them better fit the underlying data distribution, which also has downstream benefits (Mohoney et al., 2024).
Various variance measures exist for detecting imbalances in cluster assignments (Hastie et al., 2009). We selected the following subset as they capture different aspects of imbalance, both locally and globally.
RefineHeuristic: A local region is considered a good candidate for refinement (returning True) if any of the following conditions hold:
Refine: Rule 1 captures local imbalance using the coefficient of variation:
[…]
@@
Refine: Rule 2 captures extreme values using local spread:
[…]
This ensures that refinement is triggered when large disparities exist between the smallest and largest clusters within a local region.
Refine: Rule 3 captures imbalance based on the global distribution:
[…]
where $c_{1},c_{2}$ are cracks in the local region, and $S(c)=H_{C}^{\text{true}}[c]$ represents the size (i.e., number of assignments) of crack $c$.
[…]
These thresholds ensure that refinement is applied only when clear imbalance exists, maintaining index efficiency. An optimized implementation could relax these rules when more computational budget is available, allowing refinement of less imbalanced regions.

### v2

@@ 3.3. Outline of the solution
We propose a partition-based crackable IVF index for ANN search. It is designed to remain near the minimum cumulative time across the number of queries submitted, be constructed as the query load increases, and avoid indexing unused parts of the vector space.
At initialization, CrackIVF starts as a small IVF index, with coarse partitioning of the entire dataset. As queries come in, more partitions are added in the accessed regions related to the queries (Figure 2(c)). The index employs two localized and independent build operations: CRACK and REFINE. The former introduces new partitions, while the latter applies a localized K-means to refine them. These operations run independently across both time and region of the vector space. Queries and their local visited regions act as CRACK and REFINE candidates, determining where new partitions should be introduced. Using the queries to create partitions, rather than selecting random ones, allows the distances computed during search to be used for stealing points from the local region, that are closer to the query than their current assignment. Thus, intermediate results from search operations are reused to amortize the cost of construction. Additionally, using the queries as candidates for CRACK and REFINE enables CrackIVF to follow the query distribution in terms of when and where to execute build operations. Partitions that receive a larger number of queries will be proportionally more likely to be selected for cracking and refinement.
Finally, to mitigate the high overhead of storage reorganization in high-dimensional IVF indexes, we do not perform build operations after every query. Instead, CRACK operations are buffered and committed only when it is beneficial, while REFINE operations are executed eagerly, but infrequently, to improve local regions that queries are visiting. We introduce two control mechanisms to make these decisions. The first is a set of heuristic rules that determine the decision boundaries for where in the vector space cracks should be buffered or refinements executed. The second is a cost-based approach that estimates the latency overhead of a build operation and constrains it relative to the cumulative runtime of the index since initialization. As more queries arrive and the index is used more frequently, the budget for build operations increases proportionally, allowing the index to improve as the query workload demands increase.
Table 1. Notation used in algorithms and cost model.
@@
$metric(a,b)$ Similarity metric, where in our work
$metric\in\{\text{Eucl. Distance (L2)},\text{Inner Product (IP)}\}$
@@
$C_{\text{local}}\in\mathbb{Z}^{nprobe}$ Local region crack IDs
$P_{\text{local}}$ Local region points
@@
$\max_{\text{points}}$ K-means max training sample per crack
nlistnlist Total inverted lists storing $P$, where $nlist=|C|$
@@
bsbs Total queries batched in single SEARCH
nprobenprobe Number of partitions scanned per SEARCH
@@
$\alpha$ Ratio of total build time to total time
$A_{P}^{\ast}\in\mathbb{Z}^{|P|}$ Point assignments to crack IDs
$D_{P}^{\ast}\in\mathbb{R}^{|P|}$ Point distances to crack representative
$H_{C}^{\ast}\in\mathbb{N}^{|C|}$ Histogram of points per crack
Note: $\ast\in\{\text{dyn},\text{true}\}$, where dyn refers to the dynamic state (includes buffered cracks) and true refers to the true index state.
@@ 4. Crack-IVF
CrackIVF, is implemented in FAISS (Douze et al., 2024). We describe the core logic of the algorithm and operations in Section 4.1 and Algorithm 1; the control mechanisms, including heuristic rules and cost models, in Sections 4.2 and 4.3. These aid in making accurate decisions about where and when to perform the build operations, CRACK and REFINE (Algorithms 2 and 3). The notation used is in Table 1.
@@ 4.1. Algorithm and Operations
CrackIVF performs index construction and storage reorganization operations as a side-effect of query execution to incrementally improve the index. It has three core operations, $SEARCH$ finds the k-NN results to return to the user, while $CRACK$ and $REFINE$ operations reorganize the IVF structure and change the index state.
SEARCH: During the search procedure of IVF, the local region to the query is identified by finding the nprobenprobe nearest partitions to the query. Subsequently, the point assignments to these local partitions are scanned to find the $k$ nearest neighbors. The result includes the nearest point ids $I$, their distances $D$ to the query, and the $C_{visited}$ ids of the local region cracks that the query visited.
@@
8:   $I_{\text{steal}}\leftarrow\{p\in I_{\text{local}}\mid D_{\text{local}}(p)\mathrel{\text{op}}D_{P}^{\text{dyn}}(p)\}$ {op = << for L2, >> for IP}
9:   good_crack_candidategood\_crack\_candidate = CrackHeuristic($I_{steal},\text{State}_{\text{dyn}}$)
10:   if good_crack_candidategood\_crack\_candidate then
@@
20:  if can_afford_crackcan\_afford\_crack and enough_bufferedenough\_buffered then
@@
[…]
@@
34:    if can_afford_refinecan\_afford\_refine then
@@
36:     if good_refine_candidategood\_refine\_candidate then
@@
CRACK: Physical reorganization of points between inverted lists, and the mutation of the index does not happen after each query. Each incoming query represents a crack candidate, which is evaluated by rules that define the decision boundaries which classify it as good or bad (Section 4.2) (Line 9 Algorithm 1). Only good cracks are further buffered and potentially committed at a later time. Each crack is essentially a new partition, where the points assigned to the partition have been “stolen” based on their proximity to the query. We reuse the results of the SEARCH operations to compare if the distance of a point to the query is closer than the distance of the point to its current assignment. If it is, then it’s beneficial to reassign the point to the new crack, and this assignment is buffered, to be committed when $CRACK$ is executed, as explained in Section 4.3. $CRACK$ is the procedure that performs the physical reorganization, synchronizing the buffered state of the index with the currently working state. The core individual kernels of the procedure are shown in Algorithm 2. It adds a new inverted list for each buffered crack, moves stolen points to their new assignments, and updates the partition representatives to be the centroid of the points based on their new assignments. This procedure is only executed when enough budget for build operations has been accumulated throughout the usage of the index (Line 18 Algorithm 1). This budget control mechanism ensures that CrackIVF does not spend a disproportionate amount of time on CRACK operations. The parameter $\alpha$, the ratio of total build operations time to the total time, is what controls the available budget. Since $CRACK$ is applied lazily, the heuristics governing where a crack is added and the cost-based control mechanism for when it is added are independent.
@@
REFINE: Refinement is the second operation that improves the index. It is performed eagerly, focusing on the local region visited by the query. A local region is defined by the nprobenprobe nearest partitions that the query visits during search. Algorithm 1 Line 7, is where the local region $C_{local}$ for a query is extracted from the SEARCH result $C_{visited}$ of a query batch. The goal of REFINE is to immediately correct suboptimal assignments in the regions of the current index. Thus, any buffered cracks and assignments within the local region are not considered, as they are not part of the current index state. The decision to refine a region is based on heuristics evaluating the imbalance of point assignments in local cracks. Executing a refinement operation depends on an estimate of the runtime cost, determined using a cost model specific to $REFINE$ (Line 33, Algorithm 1). Unlike cracking, the opportunity to refine a local region takes place when both the currently visited local region is a good candidate for refinement and there is sufficient computational budget to execute the operation. If the “where” and “when” heuristics for $REFINE$ do not simultaneously agree, the next opportunity is when another query accesses the same region. $REFINE$ is shown in Algorithm 3). It executes a sequence of kernels implementing a local K-means variation and performs physical reorganization. This process improves the placement of crack representatives within the local region and reassigns points to their nearest representative. $REFINE$ has higher computational costs as it does not rely on precomputed buffered assignments, and assignments are determined dynamically upon completion of the local K-means process. $CRACK$ has a higher data-movement cost, since typically a larger region of the vector space is affected when cracks are added throughout.
@@
2:  Local K-Means: Refine region of $C_{\text{local}}$, producing $C^{\prime}_{\text{local}}$.
@@
Index State: Since CRACK is a lazy operation, the index has two concurrent states, the true current state and a dynamic buffered state. The true state reflects the current structure of the index, while the dynamic one tracks its expected future structure for when $CRACK$ is executed, which synchronizes the two states. $REFINE$ operations directly alter the current index state, while cracking decisions are made based on the dynamic state. This can lead to scenarios that require us to undo buffering decisions. This is done to avoid bad cracks from being committed and having a situation where we commit decisions to the index that we know to be suboptimal (Line 21, Algorithm 1). To track the index state, we define:
(1) $\text{State}_{\ast}=(A_{P}^{\ast},D_{P}^{\ast},H_{C}^{\ast})$
@@
where $\ast\in\{\text{dyn},\text{true}\}$ denotes either the dynamic uncommitted or true committed state. The dynamic state consists of dynamic assignments $A_{P}^{\text{dyn}}$, dynamic point distances $D_{P}^{\text{dyn}}$, and a histogram $H_{C}^{\text{dyn}}$ of both the true and buffered crack sizes. Similarly, the true state represents the true, committed state of assignments, distances, and crack sizes in the index. We extend the FAISS-IVF “.add()” operation to return the assignments and distances computed, which are an intermediate result, in order to initialize the state tracking. The state held in $A_{P}^{\text{dyn}}$$D_{P}^{\text{dyn}}$ and $H_{C}^{\text{dyn}}$, is only consistent with the actual state of the index after $CRACK$ operations and before buffering any cracks. If no cracks are buffered and only $REFINE$ operations are executed, the dynamic and true state are always consistent. All of the above state metadata is used after the SEARCH operation, along with the search results $I,D,C_{visited}$ to classify each query as a good or bad crack candidate (Section 4.2).
@@
Rollback of buffered state: Consider the following scenario. A $CRACK$ operation was just executed at time $t_{0}$, so $State_{true}=State_{dyn}$. After a few queries, at $t_{1}$, assume that 2 good crack candidates have been buffered, both in the same $C_{local}$ region of the vector space and thus now $State_{true}\neq State_{dyn}$. At a later point, $t_{2}$, a $REFINE$ is executed, also in $C_{local}$ changing the assignments to $C^{\prime}_{local}$, affecting $S_{true}$. Since $REFINE$ improves the current state eagerly, the decision that the buffered cracks are good was taken at a prior time $t_{1}$, on a prior state $C_{local}$. Thus, at $t_{2}$, during $REFINE$ the buffered cracks may now have had their points stolen back, and assigned to other refined cracks in $C^{\prime}_{local}$. Now, while still $|C_{buffered}|=2$, it may now contain bad cracks. Finally, when at $t_{3}$, it is a good time to execute a $CRACK$ operation again, all buffered cracks are re-evaluated, as shown in lines 22-24 Algorithm 1, and only the good cracks that remain are committed.
@@
Memory Overhead and Optimizations: The additional memory cost of maintaining $State_{\text{dyn}}$ and $State_{\text{true}}$ is $O(4|P|+2|C|)$. It is negligible since $d\gg 4$ and the in-memory data alone are $O(|P|\times d)$. Additionally, these array structures are freed once the index converges or incoming queries stop. In our case, it enables an efficient compute-storage trade-off, deferring cracking costs instead of paying them per query. There are many optimizations that a full-fledged implementation could utilize. For embedding lakes, where vectors are stored on disk, state tracking arrays can be stored alongside data points instead of in the index (irrelevant for our in-memory FAISS implementation). Since queries for an individual index are typically infrequent, $State_{\text{dyn}}$ can be committed or flushed between queries, while $State_{\text{true}}$ recomputed as needed. Full memory overhead is only necessary during bursty query loads with limited hardware resources, where cracking must be buffered to reduce cumulative overhead and prevent long tail response times. Finally, the cracking overhead can be hidden from query response times since results are available before any cracking operations begin. A separate thread pool can handle cracking operations while queries use the current index state, provided $CRACK$ and $REFINE$ remain atomic to prevent inconsistencies. Prior work on the related problem of index maintenance (Xu et al., 2023) demonstrated how to concurrently execute queries by dispatching such tasks to background threads. In our work, the entire thread pool is dedicated either to search or cracking operations. We also assume a continuous query load at system capacity. As a result, the memory overhead for index state tracking is unavoidable. We further cover index maintenance prior work in Section 6.
@@ 4.2. Where to apply build operations?
Each arriving query is a crack candidate, and the visited region during index traversal is a candidate for refinement. This results in a continuous stream of candidates of varying quality. As discussed, not all vector space regions are equally important, and not all operations lead to performance gains. The goal is to maintain a stream of high-quality candidates, classified from incoming queries, enabling the index to grow by adding cracks in frequently accessed regions that need further partitioning, and refining poorly clustered regions with K-means. CrackIVF employs two heuristics, acting as a binary classification model for candidates. Each uses a set of decision rules that are distilled from our empirical observations of IVF index behavior. We do not claim to have defined an exhaustive list of decision rules for the model. Changing or learning the model and the parameters while still achieving the stated goal above, would not change the overall design of the index (Algorithm 1).
Where to CRACK (CrackHeuristic): We observe that the optimal performance is achieved at some maximum number of partitions (Figure 2(a)). Thus, for it to grow, we treat each incoming query as a good crack candidate by default and define a set of decision rules to classify when a query is a bad candidate. Query candidates classified as good, are buffered until a future $CRACK$ operation. A bad candidate (returning False) is if any of the following hold:
Don’t Crack: Rule 1 rejects queries that steal too few points:
(2) $|I_{\text{steal}}|<\text{min\_pts}$
Don’t Crack: Rule 2 prevents excessive partitioning, constraining local partition count based on the local number of points:
@@
We set $\text{min\_pts}=2$ in order to avoid adding empty cracks, and $\text{pts\_crack\_thr}=64$ to ensure an average of at least 64 points per crack locally. Rule 2 is especially important in high skew datasets, where most crack candidates are for the same region.
Where to REFINE (RefineHeuristic): We observe that refining an index improves performance at each index size (Figure 2(a)), by better distributing points across partitions and repositioning centroids to better fit the data. Optimal performance is reached at the maximum index size, so we want to avoid getting stuck refining a small index. Unlike $CRACK$, where latency is amortized by buffering, $REFINE$ is eagerly applied to immediately improve the current index, but it is also a costly operation. Thus, we treat each incoming refine candidate as bad, leaving more of the cost budget for cracks so that the index can grow. A refine candidate is the local visited region of the query, visited during search. It is a good candidate for refinement (returning True) if any of the following hold:
Refine: Rule 1 captures local imbalance:
(4) $\text{CV}_{\text{local}}=\frac{\sigma_{\text{local}}}{\mu_{\text{local}}}>\text{cv\_max}$
@@
Refine: Rule 2 captures imbalance using the global distribution:
(5) $\exists c_{1},c_{2}\quad\text{s.t.}\quad S(c_{1})\leq{\text{cutoff}_{\text{low}}},$
$S(c_{2})\geq{\text{cutoff}_{\text{high}}}$
where $c_{1},c_{2}$ are cracks in the local region, and $S(c)=H_{C}^{\text{true}}[c]$ represents the size (i.e., number of assignments) of crack $c$. The $\text{cutoff}_{\text{low}}$ and $\text{cutoff}_{\text{high}}$ correspond to cluster sizes at the size_prctlsize\_prctl and 100size_prctl100-size\_prctl percentiles of the global partition sizes. This rule captures the unlikely scenario that a large and small crack globally are next to each other locally, indicating points can be better distributed.
We set $\text{cv\_max}=2$, meaning refinement is triggered if cluster sizes vary by at least twice the local mean, and $\text{size\_prctl}=10$, defining small and large cracks based on the 10th and 90th percentiles of global cluster sizes.